# Chasing Anharmonicities in Polarization-Orientation Raman Spectra of Acene Crystals with Machine Learning

### v1

@@ V Conclusions
Based on the analysis we have presented in this paper, we can propose a way forward. The ability of the RGDOS-ML framework to decompose the PO-Raman patterns into contributions from different vibrational modes in a unambiguous manner could be used to filter intensities in PO-Raman maps and avoid the need to fit the entire pattern at once. More specifically, this could be done by considering the mode-resolved version of Eq. 14 and using it to assign PO-Raman intensities based on the self and cross-correlation terms that result from this analysis. This first-principles decomposition could then be used as a proxy for the analysis of the experimental PO-Raman maps through a more robust global fitting procedure that takes mode mixing and non-trivial spectral overlap into account [10].
@@
Nevertheless, our framework demonstrates that, starting from a given theoretical baseline, we can systematically include temperature and dynamical (anharmonic) effects on top of it, providing a predictive route to assess their impact on the PO-Raman response.
Overall, our results demonstrate that ML-driven simulations combined with established theoretical frameworks can unravel complex anharmonic phenomena in molecular crystals. The insights gained here advance the interpretation of PO-Raman spectra and provide a foundation for future improvements, including more robust deconvolution methods for analysis of 2-dimensional polarization orientation Raman patterns and larger-scale simulations to further bridge theory and experiment.
@@ VI.1 Training Machine Learning Models for Interatomic Interactions
In this paper we investigate two types of MLIPs based on neural-network (NN) architectures. One architecture is that of high-dimensional neural-network potentials (HDNNP) [54, 64], which uses two-body and three-body local atomic descriptors in the form of atom-centered symmetry functions. In this representation, there only a few hidden layers and a few nodes on the NN attached to each atomic species are needed, and the construction of the potential is successful for quite complex systems [65, 66, 67, 68, 69]. This architecture does not explicitly encode the equivariance of the transformation of each component of the symmetry functions. The target of the optimization of the NN parameters is the total energy of the system, constructed from the sum of the individual contributions of the atomic energies predicted by the NN, and forces are readily obtained by back-propagation (i.e., the model is conservative). We have referred to this model as HDNNP-MLIP in the main text.
@@
For naphthalene, we used the MACE potential decribed in Ref.[17]. The underlying electronic-structure is also PBE+MBD with tight basis-set. For further details on the training procedure, the reader is referred to Section II of Ref.[17]. The errors on energies and forces are reported here (Table 3) for completeness.
@@ VI.3 Phonons calculations and geometry optimizations
For the comparison of Section IV.1 we computed the harmonic phonon frequencies for anthracene on the minimum energy atomic configuration with lattice vectors fixed to the experimental structure at 295K from Ref. [8] with both our ML models and DFT. The geometry optimizations were separately performed with FHI-aims (PBE+MBD, tight basis set) for the DFT case and with our two ML potentials. Phonon calculations were subsequently performed using the phonopy code [81, 82].
VI.4 Raman tensors
The Raman tensors needed for the RGDOS expansion were obtained from finite differentiation of DFPT polarizabilities of displaced minimum energy structures along the normal coordinates. The procedure is repeated for every material at every different lattice parameters considered. The PBE functional was employed in the DFPT calculations with tight basis-sets. Again, a comparison of LDA vs PBE Raman tensors is reported in Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results..
@@ VI.6 Fitting of PO-dependence
Following Ref. [8] we fitted the PO-dependence of the integrated intensity of every peak in the PO-maps, as presented in Figs. 4 and 5 for example, to a multi-Lorentzian function of the form
@@ VI.7 VDOS and CVDOS calculations
Data pertaining to this paper will be made available in a public repository upon publication.

### v2

@@ V Conclusions
Based on the analysis we have presented in this paper, we can propose a way forward. The ability of the RGDOS-ML framework to decompose the PO-Raman patterns into self- and cross-contributions from different vibrational modes could be used to direclty assign intensities to a given spectral feature in the simulated PO-Raman maps, avoiding the need to fit the entire pattern at once. While in this manuscript we chose to compare on equal grounds the global fit of both the simulation and experiment, this alternative first-principles decomposition could be used in the future to guide the experiments towards a well-informed fitting procedure that takes mode mixing and non-trivial spectral overlap into account [12].
@@
Nevertheless, our framework demonstrates that, starting from a given theoretical baseline, we can systematically include temperature and dynamical (anharmonic) effects on top of it, providing a predictive route to assess their impact on the PO-Raman response. Overall, we showed that ML-driven simulations combined with established theoretical frameworks can unravel complex anharmonic phenomena in molecular crystals. The insights gained here advance the interpretation of PO-Raman spectra and provide a route for future improvements, including more robust deconvolution methods for analysis of 2-dimensional polarization orientation Raman patterns and larger-scale simulations to further bridge theory and experiment.
@@ VI.1 Training Machine Learning Models for Interatomic Interactions
In this paper we investigate two types of MLIPs based on neural-network (NN) architectures. One architecture is that of high-dimensional neural-network potentials (HDNNP) [10, 11], which uses two-body and three-body local atomic descriptors in the form of atom-centered symmetry functions. In this representation, there are only a few hidden layers and a few nodes on the NN attached to each atomic species are needed, and the construction of the potential is successful for quite complex systems [59, 3, 2, 71, 58]. This architecture does not explicitly encode the equivariance of the transformation of each component of the symmetry functions. The target of the optimization of the NN parameters is the total energy of the system, constructed from the sum of the individual contributions of the atomic energies predicted by the NN, and forces are readily obtained by back-propagation (i.e., the model is conservative). We have referred to this model as HDNNP-MLIP in the main text.
@@
For naphthalene, we used the MACE potential decribed in Ref.[33]. The underlying electronic-structure is also PBE+MBD with tight basis-set. For any further detail on the training procedure, the reader is referred to Ref.[33]. The errors on energies and forces are reported here (Table 3) for completeness.
@@ VI.3 Phonons calculations and geometry optimizations
For the comparison of Section S3 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. we computed the harmonic phonon frequencies for anthracene on the minimum energy atomic configuration with lattice vectors fixed to the experimental structure at 295K from Ref. [4] with both our ML models and DFT. The geometry optimizations were separately performed with FHI-aims (PBE+MBD, tight basis set) for the DFT case and with our two ML potentials. Phonon calculations were subsequently performed using the phonopy code [76, 74].
VI.4 Raman tensors for -RGDOS
The Raman tensors needed for the RGDOS expansion were obtained from finite differentiation of DFPT polarizabilities of displaced minimum energy structures along the normal coordinates. The procedure is repeated for every material at every different lattice parameters considered. The LDA functional was used for the benchmarks of Section III.3 for a direct comparison with the LDA-trained SA-GPR model, while the PBE functional was employed in the production calculations of the -RGDOS spectra. All DFPT calculations employed tight basis-sets. While our comparison of LDA vs PBE Raman tensors in Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. shows nearly identical results, we opted for PBE in production calculations due to reduced computational cost of the RGDOS methodology, and to ensure full consistency with the PBE level of theory of the MLIPs employed. Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (see also additional Ref. [75] therein) also includes a brief discussion on a posteriori symmetryzation of numercially noisy Raman tensors.
A further benchmark unpolarized harmonic Raman spectra calculated at different levels of theory is provided for completeness in Section S13 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. We show how different DFT functionals (including hybrid-GGA) and dispersion corrections do not significantly alter Raman intensities in anthracene at the harmonic level.
@@ VI.6 Fitting of PO-dependence
Following Ref. [4] we fitted the PO-dependence of the integrated intensity of every peak in the PO-maps, like the ones in Figs. 3, to a multi-Lorentzian function of the form
@@ VI.7 VDOS and CVDOS calculations
Data pertaining to this paper will be made available at https://github.com/sabia-group/Paper-PO-Raman upon publication.